On October 1, 2017 Marigold Corp. issued 5%, 10-year bonds with a face value of $7920000 at 103. Interest is paid on October 1 and April 1, with any premiums or discounts amortized on a straight-line basis. Bond interest expense reported on the December 31, 2017 income statement of Marigold Corp. would be
Answer #1

Solution

Interest expense on December 31 $        93,060.00

Working

Bond issue price (7920000/100*103) $ 8,157,600
Face value $ 7,920,000
Premium on bonds payable $ 237,600
Number of Interest payments (10 years x 2)                          20
Discount/ premium to be amortized per Half year $ 11,880
Cash Interest on bond (7920000 x 2.5%) $ 198,000
Interest expense to be recorded semiannual (198000-11880) $ 186,120
Interest expense for 3 months (186120/2) $ 93,060
Answer #2

SOLUTION :


Face value of the bond = 7920000


Premium amount = 7920000 * (103 - 100)/100 = 237600


Maturity period of the bond = 10 years.


Coupon rate is 5%, paid semi-annually

Hence, coupon amount per period of payments

= Interest expense on the bond per period

= 7920000 * 0.05/2 

= 198000


Premium of 2376000 is to be amortised for 20 semi-annual periods


As per straight line method amortisation per period

= 2376000 / 20 

= 118800


So, total expense (considered as effective interest) per period

= Coupon amount - amortised premium

= 198000 -  118800

= 79200.



Bond  issued on October 1, 2018, so effective interest for 3 months is applicable in the year ending on Dec. 31, 2017.


Hence, effective interest expense in 2017 :


= 1/2 * interest per semi-annual period

= 1/2 * 79200 

= 39600 (ANSWER)

answered by: Tulsiram Garg

> There are calculation errors. Please see the correct answer that follows it.

Tulsiram Garg Sat, Dec 4, 2021 4:44 AM
Answer #3

SOLUTION : CORRECT ANSWER :


Face value of the bond = 7920000


Premium amount = 7920000 * (103 - 100)/100 = 237600


Maturity period of the bond = 10 years.


Coupon rate is 5%, paid semi-annually

Hence, coupon amount per period of payments

= Interest expense on the bond per period

= 7920000 * 0.05/2 

= 198000


Premium of 237600 is to be amortised for 20 semi-annual periods


As per straight line method amortisation per period

= 237600 / 20 

= 11880


So, total expense (considered as effective interest) per period

= Coupon amount - amortised premium

= 198000 -  11880

= 186120



Bond  issued on October 1, 2018, so effective interest for 3 months is applicable in the year ending on Dec. 31, 2017.


Hence, effective interest expense in 2017 :


= 1/2 * interest per semi-annual period

= 1/2 * 186120 

= 93060 (ANSWER)

answered by: Tulsiram Garg